OpenStudy (anonymous):

I keep getting different answers. I think my order of operations is wrong please help.

OpenStudy (anonymous):

with?

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). I've gotten crazy answers.

OpenStudy (anonymous):

lets do it nice an slow

OpenStudy (anonymous):

first we need \[f(2)=2^2-6\times 2-7\] what do you get?

OpenStudy (anonymous):

\[f(g(x))= 8 (x²-6x-7)\]

OpenStudy (anonymous):

no it does not mean multiply

OpenStudy (anonymous):

i made a mistake it is \[f(g(2))\] so first we need \[g(2)=2^2-6\times 2-7\]

OpenStudy (anonymous):

it is just a number what number do you get?

OpenStudy (anonymous):

what happened to the 8? we need f still

OpenStudy (anonymous):

one step at a time dear, nice an slow we will get to \(f\) in a minute

OpenStudy (anonymous):

My class said to compose the equations together like f(g(x))=f(x²2-6x-7) f(g(x))= 8 (x²-6x-7) and then simplify with the x replaced with 2 which should give you the other coordinate for the function.

OpenStudy (anonymous):

but \(f(x)=x+8\) right? not \(f(x)=8x\)

OpenStudy (anonymous):

yes? is there something I don't understand?

OpenStudy (anonymous):

yes \(f\) says to ADD \(8\) to the input, not MULTIPLY it by \(8\)

OpenStudy (anonymous):

\[f(g(x))=f(x^2-6x-7)=(x^2-6x+7)-8\] in other words

OpenStudy (anonymous):

ok that was wrong \[f(g(x))=f(x^2-6x-7)=(x^2-6x+7)+8\]

OpenStudy (anonymous):

which you probably want to write as \[x^2-6x+1\]

OpenStudy (anonymous):

dang still wrong !!

OpenStudy (anonymous):

\[f(g(x))=f(x^2-6x-7)=(x^2-6x-7)+8=x^2-6x+1\] thats beter

OpenStudy (anonymous):

so then you would simply in the parenthesis with x substituted for 2?

OpenStudy (anonymous):

\[( x²-6x-7)+8 = (2²-6(2)+1)= 4-12+1=11\]

OpenStudy (anonymous):

... and again 11 is not answer

OpenStudy (anonymous):

oh wait that's 9! but its still not one of the answers. im so lost

OpenStudy (anonymous):

No no it's -7!! ( sorry i'm dyscalculic) and that is one of the answers!!